Grade 12th passAlgebra

Given the two valid equations, where a and b are constants,

2ax+4by=2 and 4ax+16by=4.

Solve for y in terms of b only.

y=

Trisha
8 Years agoGrade 12th pass

1 Answer

Satyam Negi
8 Years ago
According to the question..
 
2ax + 4by = -2
4ax + 16by = 4
 
Multipyling the 1st equation by 2 we get..
2(2ax + 4by = -2)
= 4ax + 8by = -4
 
Now, subtracting the 1st equation by the 2nd equation, we get...
(4ax + 16by = 4) - (4ax + 8by = -4)
4ax + 16by - 4ax - 8by = 4 - (-4)
8by = 8
by=1
y = (1/b)